OpenStudy (anonymous):

A city’s current population is 500,000 people. It is growing at a rate of 2.5% per year. The equation p=500,00(1.025)^x models the city’s population growth where x is the number of years from the current year. In approximately how many years will the population be 640,000?

OpenStudy (anonymous):

substitute p= 640,000 in the given eq

OpenStudy (anonymous):

\[p=500,00(1.025)^x \rightarrow 640,000=500,00(1.025)^x\] \[\frac{640,000}{500,00}=(1.025)^x \rightarrow \frac{64}{5}=(1.025)^x\] \[12.8=(1.025)^x \rightarrow \frac{128}{10}=(\frac{1025}{1000})^x\] \[\frac{2^7}{10}=(\frac{41}{40})^x\] solve it further for x

OpenStudy (anonymous):

would it be 20 years? @dpasingh

OpenStudy (ash2326):

how did you solve this?

OpenStudy (ash2326):

@jessy1289

OpenStudy (anonymous):

is it 20 years ? @ash2326

OpenStudy (ash2326):

@jessy1289 did you use a calculator or logarithm table?

OpenStudy (anonymous):

no

OpenStudy (ash2326):

then how did you solve?

OpenStudy (ash2326):

We have \[\frac{128}{10}=(\frac{41}{40})^x\] We would need to use logarithms here. Have you studied that? @jessy1289

OpenStudy (anonymous):

no but im guessing its 20 years

OpenStudy (ash2326):

That seems like a cool guess, but if you solve it you'd be surprised.

OpenStudy (ash2326):

Just google this log(128/10) / log( 41/40)

OpenStudy (anonymous):

103? @ash2326

OpenStudy (ash2326):

right